I know the answer (36; 37; 38), but don't know exactly how to write it out.
Steve's, Jack's and Matt's ages are consecutive whole numbers. Steve is the youngest, Matt is the oldest. The sum of their ages is 111. Find their agest. Is this correct?
Steve = x; Jack = x+1; Matt = x=2 How do I proceed from here

Answer by rothauserc(4718)

Steve = x; Jack = x+1; Matt = x+2 How do I proceed from here?
use the following equation
x +x+1 +x+2 = 111
3x +3 = 111
divide both sides of = by 3
x+1 = 37
x=36
Steve is 36, Jack is 37, Matt is 38